OpenStudy (anonymous):

Hello guys, I need help with this math problem 3y^3-5y-6(y+2y^3)

Parth (parthkohli):

Distribute. Do you know how?

OpenStudy (anonymous):

I'm just learning.

Parth (parthkohli):

Distributive property says that: \(\Large \color{MidnightBlue}{\Rightarrow a(x \pm y \pm z) = ax \pm ay \pm az }\)

Parth (parthkohli):

\(\pm\) means plus or minus

OpenStudy (anonymous):

okay. will you look at what I've done so far and tell me where I've made mistakes?

OpenStudy (anonymous):

This is what I've done. -3y^3-5-6(y+2y^3)= -3y^3-5y-6(y-2y^3= -9y-5y-6(y-6y)= -20-12

OpenStudy (nali):

do u still need help?